tutors, help me please!!!!!thanks
2. Eight years ago, Melanie was twice as old as Roger is now. The sum of their present ages is no more than 41. At most, how old is Roger now?

Found 2 solutions by gsmani_iyer, AnlytcPhil:
Answer by gsmani_iyer(201)


Let Reger's age now = x

So Melanie's age 8 yrs. ago = 2x ....... (given)

Melanie's present age = 2x+8

the sum of their present age = x+2x+8 =< 41

So Let us take 41 as the sum now.

3x+8 = 41. so 3x = 41-8 = 33

x = 11.

So Roger's present age is = 11. Answer.

Answer by AnlytcPhil(1806)

Gsmani's solution is OK, but he changed the inequality to
an equal sign.  I think you should keep the inequality
symbol < all the way to the end: 

  Let Reger's age now = x
  
  So Melanie's age 8 yrs. ago = 2x   .......  (given)
  
  Melanie's present age =  2x+8
  
  the sum of their present age = x+2x+8 < 41
 
  3x+8 < 41.   so 
    3x < 41-8 
    3x < 33
     x < 11. 
  
  So Roger's present age is no more than ll.

Edwin